UT 09 Nov 2021 - Split Night with PEPSI/PETS - D/E



Summary:

  • Most of the MODS calibrations remaining from previous night completed
  • LUCI1 & LUCI2 PFSalignment
  • High SNR images of pupils taken as there was lots of light from open vent doors
  • AGw2 entry in oacontrol.conf updated with newest transform
  • previous version was saved off to oacontrol.save20211109.conf first
  • transform_collect on StoneO180pm10_luci
  • central source does not appear to be stable on mask
  • central source quite bright, mostly saturated but still can centroid
  • transform_collect on Stone_P1 using FeII filter and [2x5.0s] exposures
  • central source again not very stable in position
  • Photom. Standard with bino LUCI
  • Only time for S893-D. Why does the script take so long to run? 35 minutes for 1 standard!
  • Script does 4 x [~25 x 2.51s] in each filter
  • Switch to PEPSI-PETs a little early (03:23 UT)
Night Info:
  • D/E until 04hr UT (21hr local time)
  • Reconfigure from LUCI 20:45hr, Reconfigure from LBCs 20:30hr
  • Sunset 17:29
  • Twilight Ends 18:48
  • LST@Twilight 21:38
  • Moon set 20:07
  • Twilight Begins 05:19
  • Sunrise 06:39

Night Plan:

  • Configure for LUCI-Bino, authorize LUCI-2
  • LUCI-2 AGW Transform collection verification 1-1.5hr
    • Instructions for updating transform:
    • ssh telescope@obs1 ssh oac@oac cd /lbt/oacontrol/current/etc/ file is oacontrol.conf organized by AGW#. See the related instructions in BGTechProcedures edit file in the AGW2 section, rdwrconfig -u2 then stop and restart GCSR
    • Collect Transform Wiki Reference
    • Setup: N30FieldStop, N375 Camera, H-band filter ([FeII] if star is too bright), exposure time ~4sec DIT
    • could do LUCI-1 in parallel for free, but LUCI-2 only is less distracting in case LUCI-1 has issues
    • Conditions:
      • Thin clouds OK
      • Seeing 1.2" or better
    • List of Visible Fields:
      • Stone_O180pm10_luci.list - twilight-21hr
      • Stone_P1.list - visible twlight-21hr
      • Stone_A3_LUCI.list - visible 19hr-21hr
  • If time remains AFTER the LUCI-2 AGW transform collection & clear (any seeing)
    • Photometric Standards - LUCI - Requires clear skies - any seeing
      • Script Location: /home/lbto/Service_Observing/2021B_D_E/LUCI/PhotStd
        • OT file: LUCI_2021_Restart_V1.63.xml
        • Scripts have already been generated from OT
      • Cycles through z,J,H,K,Ks
      • Sample at least 2 airmasses - repeat stars if needed
        • Targets
        • S893-D - twilight-20:45hr
        • P247U - above 30deg 19:45-20:45hr
  • if < 30 min remains AFTER LUCI-2 AGW transform collection - reconfigure for PEPSI to allow for pointing check and get on target a bit earlier.
  • If > 30 min remains AFTER LUCI-2 AGW transform collection & any clouds present (any seeing acceptable)
    • Local Mergers - NGC 7585 Bino-LUCI (or mono)
      • any conditions or seeing
      • longer dwell time allows for measuring rotation curves at ever increasing radii
      • Please observe the telluric first (in case of early closure, or problems, for example)
      • File Location: /home/lbto/Service_Observing/2021B_D_E/PoorWeather/LocalMergers
        • OT: COBandhead_Mergers_V22.xml
        • Scripts already generated: /home/lbto/Service_Observing/2021B_D_E/PoorWeather/LocalMergers/LUCI/
        • Finder Chart (correctly oriented): NGC7585_finder.jpeg
  • If conditions are too poor for LUCI-2 Transform and skies are clear (any seeing)
    • Photometric Standards - LUCI
      • Script Location: /home/lbto/Service_Observing/2021B_D_E/LUCI/PhotStd
        • OT file: LUCI_2021_Restart_V1.63.xml
        • Scripts have already been generated from OT
      • Cycles through z,J,H,K,Ks
      • Sample at least 2-3 airmasses - repeat stars if needed
      • Targets
        • S893-D - twilight-20:45hr
        • P247U - above 30deg 19:45-204:45hr
  • If conditions are too poor for LUCI-2 transform AND skies are not clear (clouds and any seeing)
    • Local Mergers - NGC 7585 Bino-LUCI (or mono)
      • any conditions or seeing
      • longer dwell time allows for measuring rotation curves at ever increasing radii
      • Please observe the telluric first (in case of early closure, or problems, for example)
      • File Location: /home/lbto/Service_Observing/2021B_D_E/PoorWeather/LocalMergers
        • OT: COBandhead_Mergers_V22.xml
        • Scripts already generated: /home/lbto/Service_Observing/2021B_D_E/PoorWeather/LocalMergers/LUCI/
        • Finder Chart (correctly oriented): NGC7585_finder.jpeg
  • Possible MODS tasks if conditions deteriorate too badly for transform after start of night (O. Kuhn)
  • If known before sunset that LUCI-2 transform cannot be attempted:
    • LBC/TMS testing
      • TMS Setup
        • ssh telescope@robs
        • cd /lbt/lbto/supportscripts/MetrologySupport
        • Check if reference file is copied over (J. Hill or O. Kuhn will verify)
        • Go to co-point field, run dofpia
        • AFTER dofpia is completed then run TMS
          • First Set reference: ./lbc_set_tms_offset.py -s 2
            • this will take a while before returning prompt
          • In MetrologySupport directory execute ./tms_loop.pl once we are configured and authorized
        • co-point with lbcrangebal, then go to field of choice
        • Start running scripts
        • CTRL-C before running DOFPIA again - if run DOFPIA and TMS simultaneously, the two will "fight" - TMS will backout anything DOFPIA does!
    • Reconfigure at 20:30hr
### Cut-and-paste of Dave's logs:

### ECD+PEPSI - UT 20211109

luciX.20211109.NNNN.fits

LUCI1 LUCI2 Comments

0001 ---- 0001 ---- 1x0.0s dark test exposure to check RTD display

PFSalign LUCI2

0002 FS @ 1018.5 1022.0

0003 Pupil

plug @ -10 dX +20 dY

stop @

PFSalign 5.5 2.0 10.0 20.0 Cam=N375

FM1m1 FM1m2 FM4m1 FM4m2

  1. 1384 1154 -1795

0004 Pupil

-20 dX still (went wrong way above

PFSalign 0.0 0.0 -20.0 0.0 Cam=N375

FM1m1 FM1m2 FM4m1 FM4m2

-41 -2405 -3 1622

0005 Pupil looks good

0006 FS @ 1025 1024

...good enough!

0007 dark

0008 20x0.0s dark

0009 pupil

0010 FS

LUCI1

0002 FS @ 1023.4 1023.3

0003 Pupil

plug @ -10 dX +3 dY

stop @

PFSalign 0.6 0.7 -10.0 3.0 Cam=N375 ver-

FM1m1 FM1m2 FM4m1 FM4m2

531 -1197 207 654

0004 Pupil ...looks good

0005 20x0.0s LIR/INT high SNR image of pupil aligned

0006 FS @ 1023.2 1023.8

...good enough!

0007 20x0.0s dark (can still see the FS faintly)

0008 20x0.0s FS

Update AGw2 transform: note: active transform is dated 20170905!

Existing oacontrol.conf copied to oacontrol.save20211109.conf

Installed 20211109 and previous model commented out in oacontrol.conf

##########################################

#

# Measured 20211031

# PCS to AGw transformation matrix

# X coefficients

sfpagw11=-1.0069509121332778090

sfpagw12=0.0122493892772430682

X_0=-0.0849

# Y coefficients

sfpagw21=-0.0048312098527816816

sfpagw22=-1.0008980429616880503

Y_0=612.2977

# Pinhole offset in polar coordinates

# Radial (mm)

r_gpoff=0.000

# Angle (rad)

psi_gpoff=0.000

#

##########################################

Steve restarting GCSR

01:04 transform_collect using Stone_O180pm10_luci.list

start faint through FeII filter, ^C and switch to H

01:10 transform_collect using Stone_O180pm10_luci.list

on axis source is neither well-centered nor stable from one off-axis

guide star to the next.

/lbt/data/new/luci2.20211109.0011.fits

/lbt/data/new/luci2.20211109.0012.fits

/lbt/data/new/luci2.20211109.0013.fits

/lbt/data/new/luci2.20211109.0014.fits

/lbt

imdel sky

imdel luci2*

imcopy /lbt/data/new/luci2.20211109.0014 sky

imar /lbt/data/new/luci2.20211109.0015 - sky luci20015

imar /lbt/data/new/luci2.20211109.0016 - sky luci20016

imar /lbt/data/new/luci2.20211109.0017 - sky luci20017

imar /lbt/data/new/luci2.20211109.0018 - sky luci20018

imar /lbt/data/new/luci2.20211109.0019 - sky luci20019

imar /lbt/data/new/luci2.20211109.0020 - sky luci20020

imar /lbt/data/new/luci2.20211109.0021 - sky luci20021

imar /lbt/data/new/luci2.20211109.0022 - sky luci20022

imar /lbt/data/new/luci2.20211109.0023 - sky luci20023

imar /lbt/data/new/luci2.20211109.0024 - sky luci20024

imar /lbt/data/new/luci2.20211109.0025 - sky luci20025

imar /lbt/data/new/luci2.20211109.0026 - sky luci20026

imar /lbt/data/new/luci2.20211109.0027 - sky luci20027

imdel erme

imcombine luci2* erme combine=sub reject=none

02:08 Restart transform_collect on Stone_P1 using 2x5.0s in FeII

02:45 Switch to Bino LUCIs

Photometric Standards - LUCI - Requires clear skies - any seeing

Script Location: /home/lbto/Service_Observing/2021B_D_E/LUCI/PhotStd

OT file: LUCI_2021_Restart_V1.63.xml

02:47 S893-D - twilight-20:45hr

finished at 03:23

Script does 4 dithers on each filter, 4 x [~1 min] exposures

imdel erme*.fits

imar /lbt/data/new/luci2.20211109.0073 - /lbt/data/new/luci2.20211109.0074 erme_z1

imar /lbt/data/new/luci2.20211109.0075 - /lbt/data/new/luci2.20211109.0076 erme_z2

imar /lbt/data/new/luci2.20211109.0077 - /lbt/data/new/luci2.20211109.0078 erme_Ks1

imar /lbt/data/new/luci2.20211109.0079 - /lbt/data/new/luci2.20211109.0080 erme_Ks2

imar /lbt/data/new/luci2.20211109.0081 - /lbt/data/new/luci2.20211109.0082 erme_K1

imar /lbt/data/new/luci2.20211109.0083 - /lbt/data/new/luci2.20211109.0084 erme_K2

imar /lbt/data/new/luci2.20211109.0085 - /lbt/data/new/luci2.20211109.0086 erme_H1

imar /lbt/data/new/luci2.20211109.0087 - /lbt/data/new/luci2.20211109.0088 erme_H2

imar /lbt/data/new/luci2.20211109.0089 - /lbt/data/new/luci2.20211109.0090 erme_J1

imar /lbt/data/new/luci2.20211109.0091 - /lbt/data/new/luci2.20211109.0092 erme_J2

display erme_z1 1 zsc- zra+ ; imexam ## 2.102E7 2.151E7

display erme_z2 1 zsc- zra+ ; imexam ## 2.221E7 2.080E7

display erme_Ks1 1 zsc- zra+ ; imexam ## 2.324E7 2.311E7

display erme_Ks2 1 zsc- zra+ ; imexam ## 2.318E7 2.297E7

display erme_K1 1 zsc- zra+ ; imexam ## 3.011E7 2.982E7

display erme_K2 1 zsc- zra+ ; imexam ## 3.030E7 3.054E7

display erme_H1 1 zsc- zra+ ; imexam ## 3.923E7 3.758E7

display erme_H2 1 zsc- zra+ ; imexam ## 3.995E7 3.872E7

display erme_J1 1 zsc- zra+ ; imexam ## 4.202E7 4.245E7

display erme_J2 1 zsc- zra+ ; imexam ## 4.289E7 4.247E7

= (2.102E7 + 2.151E7 + 2.221E7 + 2.080E7) * 2.05

P247U - above 30deg 19:45-20:45hr

no time

tstat STDIN

2.102E7

2.151E7

2.221E7

2.080E7

-- BarryRothberg - 04 Nov 2021
Topic revision: r7 - 09 Nov 2021, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback